OpenStudy (anonymous):

What is the answer for 1+1 ?

OpenStudy (saifoo.khan):

\[\Huge 11\]

OpenStudy (anonymous):

11

OpenStudy (anonymous):

11 i thnk....

OpenStudy (anonymous):

Ok thanks! very much!!!!!

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[\ln \sqrt{_{9}^{11}}\]

OpenStudy (anonymous):

\[-2e^{i\pi}\]